if cos θ= -3/4 , pie/2 < θ < pie find the value of sinθ and tanθ
Answer by mananth(16949)

cos theta = -3/4
cos^2 theta = 9/16
1-sin^2 theta = 9/16
sin^2 theta = 1-9/16
sin^2 theta = 7/16
sin Theta = +/- sqrt(7)/4
but theta is in the II quadrant
sin Theta = sqrt(7)/4